You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100616 (22 Jun 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 June15-16

Observer: V. Testa, K.Boutsia
Telescope Operator: S.Allanson
Support Astronomer: O. Kuhn
Instrument: LUCIFER1 - LBC

Summary

Candidate photometric night - Standard fields were acquired.
Problems with the dithering in LBC scripts (if the 1st dither of the OB is with offset, the offset is not performed)
30min lost during LUCI MOS AQU: OSS was not responding - committing telescope hung - all offsets were lost

Data Summary

Project Objects Filter ExpTime Notes
Mannucci HDFN 210zJHK K 1hr arc- flat- tell

Ferraro n5466 B-bessel, r-SLOAN 40min  
Giallongo Q1623 Uspec, z 1hr seeing <1.0
Piotto NGC 6791 V-BESSEL, i-SLOAN 65min  
Rigliaco Serpens Uspec, I-bessel 20min out of spec, high background

Details

02:30 - Opening the dome - starting interface

02:38 - test Flats Br-gamma to monitor sky brightness 0001

02:40 - Sky Flats Ks - imaging * 2 secs / first 20 vignetted by guideprobe 0002-0051

02:53 - Slew to Pointing Star WT10_0385 0052-0053
  • Get current pointing model IE and CA:
  • Current IE = -84.5313
  • Current CA = -61.8785
  • Corrected IE = -43.3101
  • Corrected CA = -50.2313
  • Sending commands to system to update pointing model
  • ptmodify IE -43.3101
  • ptmodify CA -50.2313

02:57 - Slew to BS9145 (Persson star) for collimating

03:05 - Slew to Telluric, changed guide star, ok acquisition ok

03:08 - HIP62402 0055-0094
  • DIT=3 NDIT=10 save=normal K band probably DIT small, redo with longer DIT

03:18 - Acquisition telluric 0095
  • DIT=10 NDIT=6 save=normal 0096-0119

03:25 - MOS acquisition HDFN Mannucci =0120-0123
  • takes very long to commit offset to telescope Problems with telescope after 15min restarting TCS or similar
  • We must redo everything from scratch

04:01 - Redo acquisition from zero 0124-0126
  • 04:11 - through-slit to check alignment OK

04:13 - Mannucci_sci_K_4 HDFN 0127-0138

05:20 - Arcs and flats
  • Arcs: 0139-0142
  • Flats: 0143-0148 Redo, lamp switched off during last exposure....

05:33 - Flats 0149-0154

05:35 - SWITCH TO LBC

05:45 - New lbc software version, PC do not start. RPC Comm. failed.
  • 05:50 - Reverted to old software: OK, started, wait for telescope to be ready

06:01 - LBC switch has been completed - slewing to WT10_311 for pointing correction

06:04 - Preset interrupted because LBT hit azimuth limits. * Azimuth motor control unit shut off. * fiber link control error for the red camera. Turn LBC OFF and reboot the CMU * AZ motors have hardware problems

06:15 - Restarted the camera
  • Hear the telescope moving...hopefully we'll observe, fiber link permitting....

06:20 - Send 10Bias_Bino

06:24 - Aborted biases, slewing to WT10_311 and check pointing
  • NEW CA, IE: -72.7 -36.5
  • Corrected IE = -36.535357919225
  • Corrected CA = -72.730946706043

06:27 - SLew to NGC 5466
  • dofpia
  • Blue pupils too small Z4 -10000
  • Red strange fits Z8 +2000
  • RED 294 205 -20 -316 -476 392 -13 0.57 0.39
  • BLUE -7087 0 0 0 0 0 0 0.73 2.79
  • RED -295 13 127 -21 -65 235 6 0.62 0.21
  • BLUE -401 -406 -558 262 373 -413 -10 0.99 0.43
  • RED 27 -114 -39 130 51 33 1 0.68 0.11
  • BLUE -271 -133 63 54 35 -38 1 1.08 0.14

06:38 - NGC5466_f2_4
  • seeing ~0.7" !

06:54 - NGC5466_f2_5
  • seeing ~0.65"!

07:10 - NGC5466_field
  • airmass 1.4 but seeing very good, so we decide to execute the OB and complete the program on this object
  • 45 min without dofpia, still 3.1 pix FWHM, slight astigmatism at the bottom right corner, overall very good.

07:27 _ Slew to Q1623 - dofpia
  • RED 1208 256 -154 217 99 -350 -30 0.47 0.55
  • BLUE 3622 197 -47 67 -64 -824 -82 0.76 1.52
  • RED -284 -83 2 -23 -89 21 -14 0.45 0.15
  • BLUE -179 53 198 -30 -92 -206 -17 0.67 0.18
  • RED 146 -89 63 26 -16 -65 -2 0.47 0.11
  • BLUE 84 -66 -201 -81 67 -86 -20 0.60 0.13

07:33 - Q1623_Uspec_3
  • seeing <~ 0.7" in the U too!
  • Uspec not correcting on line AO probably because no pupils available

08:10 - dofpia
  • RED 71 36 107 59 -34 -87 -3 0.49 0.11
  • BLUE -104 -47 141 -15 -71 -439 -4 0.82 0.30
  • RED 34 -11 -167 -26 -44 -48 -4 0.47 0.10
  • BLUE 87 -49 135 74 15 -216 -8 0.77 0.17

08:14 - Q1623_Uspec_4

08:50 - SA110_BV_ri

08:58 - Slew to NGC 6791
  • dofpia
  • RED 448 -237 -159 -329 66 -41 32 0.60 0.26
  • BLUE 3992 -22 -650 244 -388 -694 -23 0.91 1.66
  • repoint on a less crowded spot
  • RED 1182 403 268 670 -88 -414 -40 0.49 0.64
  • BLUE 674 317 51 -106 375 -511 1 0.81 0.47
  • RED 83 -226 -34 -126 -64 -171 -5 0.48 0.17
  • BLUE -470 102 -289 -154 -4 -85 40 0.75 0.24
  • RED -494 -6 -118 77 2 74 -3 0.45 0.22
  • BLUE 918 -33 573 59 12 -368 -16 0.68 0.47
  • RED 292 10 188 -283 164 -91 -4 0.58 0.22
  • BLUE -172 -177 -79 -54 -155 -90 -14 0.62 0.15

09:13 - WD_N6791_LONG_03
  • hit limits in Y, incomplete , miss last dither position, some red images slightly elongated but probably still usable

09:53 - dofpia
  • RED 1077 117 -223 156 83 -271 -8 0.50 0.48
  • BLUE 1006 -93 252 190 -70 -377 38 0.86 0.48
  • RED -241 47 82 14 -17 -18 6 0.52 0.13
  • BLUE -254 311 -179 -57 -73 -73 30 0.90 0.18
  • RED 31 -99 117 -54 23 14 0 0.57 0.10
  • BLUE 36 -71 195 33 -3 -68 9 0.85 0.11

01:01 - WD_N6791_LONG_04

01:40 - Slew to Serpens and focus
  • dofpia
  • RED 721 139 -34 -244 -303 70 -10 0.65 0.35
  • BLUE -1142 -330 -585 64 247 -230 -7 1.01 0.53
  • RED not arrived stop and restart
  • dofpia
  • RED -1448 155 4 73 27 368 14 0.74 0.62
  • BLUE 13 -120 -391 96 35 -235 0 0.99 0.21
  • RED -33 -346 27 0 156 -6 4 0.71 0.15
  • BLUE 19 -163 -10 -13 -35 -47 0 1.05 0.10
  • Stopped

10:53 - Serpens UsI 1
  • Sky is already starting to get bright, try anyway

11:18 - SA110

11:26 - slew to FF * Flats done: Vi, Br, Usz

11:52 - End of night, yeah!

But before....Luci darks.

-- VincenzoTest - 16 Jun 2010
Topic revision: r4 - 22 Jun 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback